# Autoregressive Sequence Modeling for 3D Medical Image Representation

### v1

@@ Abstract
Three-dimensional (3D) medical images, such as Computed Tomography (CT) and Magnetic Resonance Imaging (MRI), are essential for clinical applications. However, the need for diverse and comprehensive representations is particularly pronounced when considering the variability across different organs, diagnostic tasks, and imaging modalities. How to effectively interpret the intricate contextual information and extract meaningful insights from these images remains an open challenge to the community. While current self-supervised learning methods have shown potential, they often consider an image as a whole thereby overlooking the extensive, complex relationships among local regions from one or multiple images. In this work, we introduce a pioneering method for learning 3D medical image representations through an autoregressive pre-training framework. Our approach sequences various 3D medical images based on spatial, contrast, and semantic correlations, treating them as interconnected visual tokens within a token sequence. By employing an autoregressive sequence modeling task, we predict the next visual token in the sequence, which allows our model to deeply understand and integrate the contextual information inherent in 3D medical images. Additionally, we implement a random startup strategy to avoid overestimating token relationships and to enhance the robustness of learning. The effectiveness of our approach is demonstrated by the superior performance over others on nine downstream tasks in public datasets.
[…]
@@ Introduction
The realm of medical imaging has witnessed significant evolution with the advent of advanced modalities such as Computed Tomography (CT) and Magnetic Resonance Imaging (MRI) (Najjar 2023). These three-dimensional (3D) medical images serve as a cornerstone for clinical diagnosis and treatment planning, providing physicians with a detailed glimpse into the inner workings of the human body (Panayides et al. 2020). Despite their widespread impact, the complexity and richness of 3D medical images present unique challenges in interpretation and analysis. The heterogeneity across different organs, the variability in diagnostic tasks, and the diversity of imaging modalities further complicate the comprehensive representation of these images (Dou et al. 2017; Kronman and Joskowicz 2016).
@@
Moreover, recent advances in self-supervised learning (SSL) have shown promising results in visual representation learning for certain tasks (Goyal et al. 2022; Oquab et al. 2023). These methods capitalize on the idea of reconstructing masked input data or contrastive learning to learn robust feature representations without the need for explicit labels. However, 3D medical images offer depth and volume, and are often sparse, posing new challenges when applying these methods to 3D medical image representations. As most existing self-supervised methods consider each image as a whole (Zhou et al. 2021b; Zhu et al. 2020; Zhou et al. 2023; El-Nouby et al. 2024), often overlooking inner and inter-correlations of 3D medical images, e.g., complex relationships among patches, modalities, and semantics. This limitation highlights the need for a more holistic and interconnected representation learning framework.
@@
3D Medical Images Analysis. Compared to 2D images, 3D medical images possess a significantly higher spatial complexity, hence conventional 2D methods are inadequate for learning 3D representation. Numerous studies have conducted meaningful explorations into 3D medical image analysis, primarily divided into methods based on 2D and 3D models (Bien et al. 2018; Ni et al. 2019; Roth et al. 2014; Yang et al. 2021). The advantage of 2D models lies in their ability to leverage vast amounts of natural images to obtain powerful pre-trained models. They can take 3D image planes as input channels, blending multiplanar data into 2D models (Moeskops et al. 2016; Prasoon et al. 2013; Roth et al. 2014), or use three adjacent slices as channels (Bien et al. 2018; Ding et al. 2017; Ni et al. 2019; Yu et al. 2018). However, these methods are incapable of learning the context of the 3D space, which is a critical aspect for accurate representation and analysis in 3D medical images.
@@
In contrast, 3D models are capable of learning more complex spatial features (Çiçek et al. 2016; Milletari et al. 2016; Roth et al. 2018). However, it is often challenging to obtain 3D models pre-trained on large-scale data. The approach proposed in this paper obtains powerful pre-trained 3D models by capturing inner-correlations and inter-correlations of 3D medical images.
@@
Self-Supervised Learning. Self-supervised learning (SSL) has emerged as a promising approach in recent years to harness unlabeled data. Initially, SSL has achieved remarkable results in the domain of natural images. The current methods are divided into two categories approximately: contrastive-based and reconstruction-based (He et al. 2020; Chen et al. 2020; He et al. 2022). The core idea of contrastive learning is to minimize the feature distance between different views of the same image while maximizing the distance between different images, thereby forcing the model to learn discriminative features for instances (Chen et al. 2020; He et al. 2020; Chen et al. 2021; Caron et al. 2021). Contrastive methods primarily focus on global representations and lack local focus, which leads to suboptimal performance in dense prediction tasks (Zhou et al. 2021a; Zhang et al. 2022). Reconstruction-based methods primarily focus on masked image modeling, where a significant portion of the image content is masked and then reconstructed (He et al. 2022; Xie et al. 2022). However, they only explore representation in 2D individual nature images without considering the unique correlation within 3D medical images.
@@
Due to the scarcity of annotated data, SSL has also garnered significant attention in the field of medical image analysis. Firstly, the researchers have designed many proxy tasks that focus on the characteristics of medical images, particularly for 3D images (Zhu et al. 2020; Zhou et al. 2021b; Tang et al. 2022). Additionally, based on the contrastive approach, researchers have proposed improvement strategies, as well as methods that integrate multiple proxy tasks (Zhou et al. 2021a, 2023). Although many designs have been proposed specifically for 3D medical images, current methods are often focused on individual images. There is a relative scarcity of approaches that utilize cross-sequence characteristics in 3D medical images for self-supervised representation learning. Taking the unique sequential correlation inherent in 3D medical images into full consideration, we propose to transform 3D images into various types of sequences. We design the autoregressive sequence modeling approach to learn inner and inter-correlation within 3D medical images.
@@
Token Sequences. As mentioned above, we first transform 3D images into patch sequences. Suppose each patch sequence has $N$ patches. The next step involves tokenizing each patch into $k$ visual tokens. All $N*k$ tokens from a patch sequence are concatenated into a token sequence $S$ following the order mentioned above as $\mbox{T}=\{\bm{t}_{1},...,\bm{t}_{Nk}\}$.
@@
The pre-training datasets are divided into several sources: individual images for spatial sequences, multimodal images for contrast sequences, and images belonging to the same semantic category for semantic sequences. For individual images, we collect 23,287 3D CT and MRI volumes from 12 public medical image datasets (RibFrac (Jin et al. 2020), TCIA Covid19 (An et al. 2020), AMOS22 (Ji et al. 2022), ISLES2022 (Hernandez Petzsche et al. 2022), AbdomenCT-1K (Ma et al. 2021), Totalsegmentator (Wasserthal et al. 2023), Verse 2020 (Sekuboyina et al. 2021), RSNA-2022-CSFD (Flanders et al. 2022), RSNA-2020-PED (Colak et al. 2021), STOTIC (Revel et al. 2021), FLARE22 (Ma et al. 2023, 2022), and FLARE23 (Ma et al. 2022)). For multimodal images, we collect 2,995 multimodal MRI scans from BraTS 2023 (LaBella et al. 2023), which is a series of challenges on brain MRI image analysis. Each scan of this dataset includes four MRI modalities (T1w, T1ce, T2w, and Flair). Images belonging to the same semantic category are obtained from the DeepLesion dataset (Yan et al. 2018), which contains 10,594 CT scans of 4,427 patients.
@@
We conducted downstream experiments in nine clinical tasks on public medical image datasets to evaluate the effectiveness of our method. These datasets cover a variety of organs, lesions, and modalities, including Task03 Liver (131 cases), Task06 Lung (64 cases), Task07 Pancreas (282 cases), Task08 Hepatic Vessel (303 cases), Task09 Spleen (41 cases), and Task10 Colon (126 cases) from Medical Segmentation Decathlon (MSD) (Antonelli et al. 2022), Left Atrium (LA) (Xiong et al. 2021) (100 cases), RICORD (Tsai et al. 2021) (330 cases) and LIDC-IDRI (Armato III et al. 2011) (1633 cases). These datasets can be categorized into 3D segmentation and 3D classification tasks. Specifically, Task03, Task09, and LA are used for organ segmentation, while Task06 and Task10 focus on tumour segmentation. Task07 and Task08 are designed for segmenting both organs and tumours. RICORD is used for COVID-19 binary classification (being COVID-19 or not). LIDC-IDRI is used for lung nodule binary classification (level 1/2 into negative class and 4/5 into positive class, ignoring the cases with malignancy level 3) similar to other researches that have used this dataset (Wu et al. 2018; Hussein et al. 2017). We randomly split the whole set into training, validation, and test at a ratio of 7:1:2 for the tasks on the MSD dataset. For the LA, RICORD, and LIDC-IDRI datasets, we follow the data split in  (Yu et al. 2019),  (Ye et al. 2024), and  (Yang et al. 2023), respectively. All the downstream datasets are unseen in the pre-training stage.
@@
Our baselines for segmentation include 1)UNETR (Hatamizadeh et al. 2022) and 2) 3D UNet (Çiçek et al. 2016), for classification include 1) ResNet (He et al. 2016) and 2) ViT-B (Dosovitskiy et al. 2020), which are trained from scratch. We further compare with three general SSL methods: 1) SimCLR (Chen et al. 2020), 2) MoCov3 (Chen et al. 2021), 3) DINO (Caron et al. 2021). The first two are based on contrastive learning. And DINO (Caron et al. 2021) introduces a self-distillation framework where a student model learns to predict the output of a teacher model. Then, we compare with three powerful methods specifically tailored for medical images: 1) PCRLv2 (Zhou et al. 2023), 2) MAE3D (Chen et al. 2023), 3) MedCoSS (Ye et al. 2024). Specifically, MAE3D (Chen et al. 2023) is a generative method, which leverages an autoencoder and mask-based pretext task to learn visual representations from medical images. PCRLv2 (Zhou et al. 2023) integrates pixel restoration and hybrid feature contrast into a multi-task optimization problem. MedCoSS (Ye et al. 2024) adopts a sequential pre-training paradigm using a continual learning approach.
@@
We first conduct comparative evaluations of the proposed approach against the SOTA methods on six CT-based segmentation tasks from the MSD challenge. As illustrated in Column2-8 of Table 1, our method demonstrates superior performance across all evaluated tasks on the MSD challenge, outperforming other advanced methods. The average Dice score of our method is 0.7088, which is notably higher than the scores achieved by SimCLR (Chen et al. 2020) and MoCov3 (Chen et al. 2021), which are 0.6448 and 0.6623, respectively. Compared with DINO (Caron et al. 2021), our method also outperforms it by a large margin. These methods are designed for general SSL, relying heavily on the negative sampling process or data augmentations, which is not well-suited for the nuanced and complex nature of medical images. We further conduct experimental comparisons with PCRLv2 (Zhou et al. 2023), MAE3D (Chen et al. 2023), and MedCoSS (Ye et al. 2024), which employ medical SSL. However, their performance gains are still limited. Our method gains an average 2.1% improvement in Dice score relative to the medical SSL methods. The improvements suggest that the proposed method can effectively leverage the contextual relationships inherent in 3D medical image data.

### v2

@@ Abstract
Three-dimensional (3D) medical images, such as Computed Tomography (CT) and Magnetic Resonance Imaging (MRI), are essential for clinical applications. However, the need for diverse and comprehensive representations is particularly pronounced when considering the variability across different organs, diagnostic tasks, and imaging modalities. How to effectively interpret the intricate contextual information and extract meaningful insights from these images remains an open challenge to the community. While current self-supervised learning methods have shown potential, they often consider an image as a whole thereby overlooking the extensive, complex relationships among local regions from one or multiple images. In this work, we introduce a pioneering method for learning 3D medical image representations through an autoregressive pre-training framework. Our approach sequences various 3D medical images based on spatial, contrast, and semantic correlations, treating them as interconnected visual tokens within a token sequence. By employing an autoregressive sequence modeling task, we predict the next visual token in the sequence, which allows our model to deeply understand and integrate the contextual information inherent in 3D medical images. Additionally, we implement a random startup strategy to avoid overestimating token relationships and to enhance the robustness of learning. The effectiveness of our approach is demonstrated by the superior performance over others on nine downstream tasks in public datasets. Code is publicly available at https://github.com/serena9525/AR-SSL4M.
@@ Introduction
The realm of medical imaging has witnessed significant evolution with the advent of advanced modalities such as Computed Tomography (CT) and Magnetic Resonance Imaging (MRI) (Najjar 2023). These three-dimensional (3D) medical images serve as a cornerstone for clinical diagnosis and treatment planning, providing physicians with a detailed glimpse into the inner workings of the human body (Panayides et al. 2020). Despite their widespread impact, the complexity and richness of 3D medical images present unique challenges in interpretation and analysis. The heterogeneity across different organs, the variability in diagnostic tasks, and the diversity of imaging modalities further complicate the comprehensive representation of these images (Zhou et al. 2023a).
@@
Moreover, recent advances in self-supervised learning (SSL) have shown promising results in visual representation learning for certain tasks (Oquab et al. 2023). These methods capitalize on the idea of reconstructing masked input data or contrastive learning to learn robust feature representations without the need for explicit labels. However, 3D medical images offer depth and volume, and are often sparse, posing new challenges when applying these methods to 3D medical image representations. As most existing self-supervised methods consider each image as a whole (Zhou et al. 2023b; El-Nouby et al. 2024; Gao et al. 2024), often overlooking inner and inter-correlations of 3D medical images, e.g., complex relationships among patches, modalities, and semantics. This limitation highlights the need for a more holistic and interconnected representation learning framework.
@@
3D Medical Images Analysis. Compared to 2D images, 3D medical images possess a significantly higher spatial complexity, hence conventional 2D methods are inadequate for learning 3D representation. Numerous studies have conducted meaningful explorations into 3D medical image analysis, primarily divided into methods based on 2D and 3D models (Ni et al. 2019; Yang et al. 2021). The advantage of 2D models lies in their ability to leverage vast amounts of natural images to obtain powerful pre-trained models. They can take 3D image planes as input channels, blending multiplanar data into 2D models (Moeskops et al. 2016; Prasoon et al. 2013), or use three adjacent slices as channels (Ding et al. 2017; Yu et al. 2018). However, these methods are incapable of learning the context of the 3D space, which is a critical aspect for accurate representation and analysis in 3D medical images.
@@
In contrast, 3D models are capable of learning more complex spatial features (Çiçek et al. 2016; Roth et al. 2018). However, it is often challenging to obtain 3D models pre-trained on large-scale data. Our proposed approach obtains powerful pre-trained 3D models by capturing inner-correlations and inter-correlations of 3D medical images.
@@
Self-Supervised Learning. Self-supervised learning (SSL) has emerged as a promising approach in recent years to harness unlabeled data. Initially, SSL has achieved remarkable results in the domain of natural images. The current methods are divided into two categories approximately: contrastive-based and reconstruction-based (Chen et al. 2020; He et al. 2022). The core idea of contrastive learning is to minimize the feature distance between different views of the same image while maximizing the distance between different images, thereby forcing the model to learn discriminative features for instances (Chen et al. 2020, 2021; Caron et al. 2021). Contrastive methods primarily focus on global representations and lack local focus, which leads to suboptimal performance in dense prediction tasks (Zhou et al. 2021a; Zhang et al. 2022). Reconstruction-based methods primarily focus on masked image modeling, where a significant portion of the image content is masked and then reconstructed (He et al. 2022; Xie et al. 2022). However, they only explore representation in 2D individual nature images without considering the unique correlation within 3D medical images.
@@
Due to the scarcity of annotated data, SSL has also garnered significant attention in the field of medical image analysis. Firstly, the researchers have designed many proxy tasks that focus on the characteristics of medical images, particularly for 3D images (Zhou et al. 2021b; Tang et al. 2022). Additionally, based on the contrastive approach, researchers have proposed improvement strategies, as well as methods that integrate multiple proxy tasks (Zhou et al. 2021a, 2023b). Although many designs have been proposed specifically for 3D medical images, current methods are often focused on individual images. There is a relative scarcity of approaches that utilize cross-sequence characteristics in 3D medical images for self-supervised representation learning. Taking the unique sequential correlation inherent in 3D medical images into full consideration, we propose to transform 3D images into various types of sequences. We design the autoregressive sequence modeling approach to learn inner and inter-correlation within 3D medical images.
@@
Token Sequences. As mentioned above, we first transform 3D images into patch sequences. Suppose each patch sequence has $N$ patches. The next step involves tokenizing each patch into $k$ visual tokens. All $N\times k$ tokens from a patch sequence are concatenated into a token sequence $S$ following the order mentioned above as $\mbox{T}=\{\bm{t}_{1},...,\bm{t}_{Nk}\}$.
@@
The pre-training datasets are divided into several sources: individual images for spatial sequences, multimodal images for contrast sequences, and images belonging to the same semantic category for semantic sequences. For individual images, we collect 23,287 3D CT and MRI volumes from 12 public medical image datasets (RibFrac (Jin et al. 2020), TCIA Covid19 (An et al. 2020), AMOS22 (Ji et al. 2022), ISLES2022 (Hernandez Petzsche et al. 2022), AbdomenCT-1K (Ma et al. 2021), Totalsegmentator (Wasserthal et al. 2023), Verse 2020 (Sekuboyina et al. 2021), RSNA-2022-CSFD (Flanders et al. 2022), RSNA-2020-PED (Colak et al. 2021), STOIC (Revel et al. 2021), FLARE22 (Ma et al. 2023), and FLARE23 (Ma et al. 2024)). For multimodal images, we collect 2,995 multimodal MRI scans from BraTS 23 (LaBella et al. 2023), which is a series of challenges on brain MRI image analysis. Each scan of this dataset includes four MRI modalities (T1w, T1ce, T2w, and Flair). Images belonging to the same semantic category are obtained from the DeepLesion dataset (Yan et al. 2018), which contains 10,594 CT scans of 4,427 patients.
@@
We conducted downstream experiments in nine clinical tasks on public medical image datasets to evaluate the effectiveness of our method. These datasets cover a variety of organs, lesions, and modalities, including Task03 Liver (131 cases), Task06 Lung (64 cases), Task07 Pancreas (282 cases), Task08 Hepatic Vessel (303 cases), Task09 Spleen (41 cases), and Task10 Colon (126 cases) from Medical Segmentation Decathlon (MSD) (Antonelli et al. 2022), Left Atrium (LA) (Xiong et al. 2021) (100 cases), RICORD (Tsai et al. 2021) (330 cases) and LIDC-IDRI (Armato III et al. 2011) (1633 cases). These datasets can be categorized into 3D segmentation and 3D classification tasks. Specifically, Task03, Task09, and LA are used for organ segmentation, while Task06 and Task10 focus on tumour segmentation. Task07 and Task08 are designed for segmenting both organs and tumours. RICORD is used for COVID-19 binary classification (being COVID-19 or not). LIDC-IDRI is used for lung nodule binary classification (level 1/2 into negative class and 4/5 into positive class, ignoring the cases with malignancy level 3) similar to other researches that have used this dataset (Wu et al. 2018). We randomly split the whole set into training, validation, and test at a ratio of 7:1:2 for the tasks on the MSD dataset. For the LA, RICORD, and LIDC-IDRI datasets, we follow the data split in  (Yu et al. 2019),  (Ye et al. 2024), and  (Yang et al. 2023), respectively.
@@
Our baselines for segmentation include 1)UNETR (Hatamizadeh et al. 2022) and 2) 3D UNet (Çiçek et al. 2016), for classification include 1) ResNet (He et al. 2016) and 2) ViT-B (Dosovitskiy et al. 2020), which are trained from scratch. We further compare with three general SSL methods: 1) SimCLR (Chen et al. 2020), 2) MoCov3 (Chen et al. 2021), 3) DINO (Caron et al. 2021). The first two are based on contrastive learning. And DINO (Caron et al. 2021) introduces a self-distillation framework where a student model learns to predict the output of a teacher model. Then, we compare with three powerful methods specifically tailored for medical images: 1) PCRLv2 (Zhou et al. 2023b), 2) MAE3D (Chen et al. 2023), 3) MedCoSS (Ye et al. 2024). Specifically, MAE3D (Chen et al. 2023) is a generative method, which leverages an autoencoder and mask-based pretext task to learn visual representations from medical images. PCRLv2 (Zhou et al. 2023b) integrates pixel restoration and hybrid feature contrast into a multi-task optimization problem. MedCoSS (Ye et al. 2024) adopts a sequential pre-training paradigm using a continual learning approach.
@@
We first conduct comparative evaluations of the proposed approach against the SOTA methods on six CT-based segmentation tasks from the MSD challenge. As illustrated in Column2-8 of Table 1, our method demonstrates superior performance across all evaluated tasks on the MSD challenge, outperforming other advanced methods. The average Dice score of our method is 0.7088, which is notably higher than the scores achieved by SimCLR (Chen et al. 2020) and MoCov3 (Chen et al. 2021), which are 0.6448 and 0.6623, respectively. Compared with DINO (Caron et al. 2021), our method also outperforms it by a large margin. These methods are designed for general SSL, relying heavily on the negative sampling process or data augmentations, which is not well-suited for the nuanced and complex nature of medical images. We further conduct experimental comparisons with PCRLv2 (Zhou et al. 2023b), MAE3D (Chen et al. 2023), and MedCoSS (Ye et al. 2024), which employ medical SSL. However, their performance gains are still limited. Our method gains an average 2.1% improvement in Dice score relative to the medical SSL methods. The improvements suggest that the proposed method can effectively leverage the contextual relationships inherent in 3D medical image data.
@@ Conclusions
Acknowledgements
This work was partially supported by Hong Kong Research Grants Council under Collaborative Research Fund (Project No. HKU C7004-22G).